OpenStudy (anonymous):

Car A and Car B are approaching each other from two places Town X and Town Y 210 km apart, running at 50 km/h and 55 km/h respectively. A) How long will it take them to meet? B) How far away is their meeting place from Town X?

OpenStudy (anonymous):

A) Let the time taken for the two cars to meet be = x hrs Then car A has travelled distance = 50x Car B has travelled distance = 55x 50x + 55x = 210 105x = 210 x = 210/105 x = 2 so the cars meet after travelling 2 hours B) Car A started from Town A and in 2 hours it would have covered a distance of 2*50 = 100 km So their meeting place is 100 km from town A